OpenStudy (anonymous):

How do you solve 3 square root of -80 + 4 square root of -125

OpenStudy (abb0t):

is it \[(-80)^\frac{ 1 }{ 3 } \] or 3√(-80) ?

OpenStudy (anonymous):

\[3\sqrt{-80}+4\sqrt{-125}\]

OpenStudy (anonymous):

\[3\sqrt{-80}+4\sqrt{-125}\]\[12 i \sqrt{5}+20 i \sqrt{5} \]\[32 i \sqrt{5} \]